OpenStudy (angel_kitty12):

Please help ASAP: A fair coin is flipped four times. Find: (A) The probability it will land up heads each time. (B) the probability it will land the same way each time (slightly different from (A))

OpenStudy (ybarrap):

A) Let p be the probability of success, in this case heads. Then probability of getting 4 heads is \(pppp=p^4\), where p=1/2. B) Here we have two cases: all heads or all tails. By symmetry, we have the same probability of four tails in a row as in Part A. So the total probability of either all heads or all tails is \(2p^4\)